40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 25 04:54:21 2013, Masayuki, Update, SUS, ETM violin mode violin1.pdfviolin2.pdf
    Reply  Sun Oct 27 03:41:06 2013, rana, Update, SUS, ETMX violin mode ETMX-Vio3-Ring.png
Message ID: 9290     Entry time: Fri Oct 25 04:54:21 2013     Reply to this: 9299
Author: Masayuki 
Type: Update 
Category: SUS 
Subject: ETM violin mode 

Summary

When PRMI + 2arms are locked yesterday, we heard the noise from suspension violin mode. For attenuation of that noise, we should design the resonant filter at that frequency and put into the ALS servo. I tried to measure the violin mode of ETMs SUS.

What I did

 1.The arms were locked by IR PDH. I used awggui to excite the suspention. I injected the Normal waveform, 10 Hz of bandwidth wave into C1:SUS-ETMs_ULCOIL_EXC. I put cheby filter in the FIlter of awggui. The order of that filter was 4, that has same bandwidth as that of injection wave and ripple was 4dB. I increase the injection gain with some ramp time(5sec). I swept from 600 Hz to 700 Hz. During that injection I saw the PDH error signal (POX11I and POY11I) in order to find resonance peak of violin mode.
 In ETMX resonances were easily found. That were at 631 Hz and 691 Hz. the 631 Hz peak was seen ALS error signal yesterday. On the other hand, I couldn't find ETMY violin mode. No peaks appeared any frequency.

2. For find the ETMY violin mode, I used dtt swept sine measurement. The excitation channel was C1:SUS-ETMs_ULCOIL_EXC. I measured the TF from excitation channel to POX11I and POY11I error signal. The measurement range was above 400 Hz and below 1000Hz,. The number of point is 600. I attached that result.
In ETMX curve, the coherence become bad near the resonant frequency of violin mode and also the TF is large. Although ETMX violin modes are obvious, ETMY violin modes are not visible. At 660 Hz, 780 Hz, 900 Hz the coherence is not good. That is because 60 Hz comb noise.

Discussion

 I attached the spectrum of the POX and POY error signal. Black and red curve is measured different time. I didn't inject any signal in both measurement, but the violin mode excitation has huge difference. Also there are peaks at beat frequency between violin mode and bounce mode(16 Hz), yaw motion(3 Hz). In ALS in-loop noise or XARM in-loop measurement, sometimes this region had big spikes. That was because of this resonance. And also that resonance peak couples to POY11I.

 I will measure the Q and design the resonant filter for ALS.

Attachment 1: violin1.pdf  19 kB  | Hide | Hide all
violin1.pdf
Attachment 2: violin2.pdf  22 kB  | Hide | Hide all
violin2.pdf
ELOG V3.1.3-